OpenStudy (anonymous):

a car of mass 1000kg travelling along a straight horizontal road accelerates uniformly from 15m/s to 25m/s in a distance of 320m. if the resistance to motion is 145N, find the driving force of the engine.

OpenStudy (anonymous):

Wrong section for the question.

OpenStudy (anonymous):

that is really a mathematics question. question from chapter mechanics that I had learnt in foundation mathematics II.

OpenStudy (anonymous):

I'm sorry man maybe I am mistaken. But you do need some formulae related to Physics here..

OpenStudy (anonymous):

its ok.... so do you know how to solve that problem

OpenStudy (cwrw238):

you use newtons 2nd law of motion Force = mass * acceleration

OpenStudy (cwrw238):

resultant force = driving force D - resisting force 145 N acceleration: speed is increased from 15 to 25 over 320 using thr formula v^2 = u^2 + 2as 25^2 = 15^2 + 640a a = 625 - 225 /640 = 400/640 = 5/8 m s^-2 using newtons law D - 145 = 1000 x 5/8 = 625 D = 770 N driving force = 770 N

OpenStudy (anonymous):

tq....

OpenStudy (anonymous):

@Cwrw: Beauty and the Brains. That's your title! Good one :)